A thin, homogenous bar of length L has insulated ends and initial temperature f(x) = x. Determine the temperature distribution u(x,t) in the bar.


1
Expert's answer
2021-03-08T18:10:47-0500

We know one dimensional equation for a rod is given as,


"\dfrac{\delta u}{\delta t} = c^2\dfrac{\delta^2 u}{\delta x^2}"


It's solution can be given as "u(x,t) = (C_1cos(px) + C_2sin(px))e^{-c^2p^2t}"


At x = 0 ,"u(x,t) = 0 = u(0,t) = C_1e^{-c^2p^2t}"


Hence, "C_1 = 0"


"u(x,t) = C_2sin(px)e^{-c^2p^2t}"


Using condition u(a,t) = 0


Hence, "C_2sin(px)e^{-c^2p^2t} = 0"


But actually it is not equal to zero to get the desired solution


then, "sin(px) = 0"


"pa = n\pi"


Hence, "p = \dfrac{n\pi}{a}"

Putting the value of p in above equation, we get


"u(x,t) = C_2sin{\dfrac{n\pi x}{a}}e^-\dfrac{n^2\pi^2kt}{a^2}"


where, n = 0,1,2,3.....


"u(x,t) = \sum_{1}^{\infin}b_nsin{\dfrac{n\pi x}{a}}e^-{\dfrac{n^2\pi^2kt}{a^2}}" .......(A)


It is given that u(x,0) = f(x)


So, "f(x) = \sum_{1}^{\infin}b_nsin{\dfrac{n\pi x}{a}}"


then "b_n = \dfrac{2}{a}\int _{0}^{a}f(x)sin{\dfrac{n\pi x}{a}}dx" (Using Fourier Sine Series)


Putting this value of "b_n" in equation (A) we get the desired equation.


"u(x,t) = \sum_{1}^{\infin} \dfrac{2}{a}\int _{0}^{a}f(x)sin{\dfrac{n\pi x}{a}}dx (sin{\dfrac{n\pi x}{a}}e^-{\dfrac{n^2\pi^2kt}{a^2}})"